Serving at a speed of 170 km/h, a tennis player hits the ball at a height of 2.5 m and an angle θ below the horizontal. The service line is 11.9 m from the net, which is 0.91 m high. What is the angle θ such that the ball just crosses the net? Will the ball land in the service box, whose out line is 6.40 m from the net?


1
Expert's answer
2021-11-15T12:35:53-0500

"170\ (km/h)=47.22\ (m/s)"


"v_{0x}=47.22\cdot\cos\theta"




"11.9=v_{0x}\cdot t_1=47.22\cdot\cos\theta\cdot t_1"


"2.5-0.91=47.22\cdot\sin \theta\cdot t_1+gt_1^2/2" . From these two equations


"0.311\tan^2\theta+11.9\tan\theta-1.279=0\to \theta=6.1°" . Answer


Time of movement in the vertical direction


"2.5=47.22\cdot\sin6.1°\cdot t_2+9.8\cdot t_2^2/2\to t_2=0.367\ (s)"


"l=47.22\cdot\cos6.1°\cdot0.367=17.2\ (m)"


The service box is placed at a distance "11.9+6.4=18.3\ (m)" . So,


the ball will not land in the service box. Answer